Laverny tosses an unbiased coin. He receives 100 if a head appears and he pays 40 if a tail appears. Find the variance of his gain.Β 


1
Expert's answer
2022-02-28T15:59:37-0500

x. 100. - 40

P(x). 0.5. 0.5


E(x)=0.5(100) +0.5(-40)

= 30

"E(x^2)=0.5(100^2)+0.5(-40)^2"

= 5800

Var(x) ="E(x^2)-E(x)^2"

= "5800-30^2"

=4900